OpenStudy (anonymous):

Given:3(x-4)=33 Prove: x=15 Need help on proofing these with the reasons in 10 steps

OpenStudy (anonymous):

Do you have to have it in at least 10 steps? I can show you in about 7 if thats ok?

OpenStudy (anonymous):

3(x-4)=33 Expand the bracket 3(x)-4(3)=33 3x-12=33 Add both sides by 12 3x=33+12 3x=45 Divide both sides by 3 to get x on its own x=45/3 x=15

OpenStudy (anonymous):

you can do it in two steps 1) divide by 3 2) add 4